OpenStudy (anonymous):

Suppose the area of a circle is 16π. What is its diameter?

OpenStudy (anonymous):

Thank you!!

OpenStudy (anonymous):

Use \[D=2 R\\Area = \pi R^2\]

OpenStudy (anonymous):

\[ 16 \pi = \pi R^2 \]

OpenStudy (anonymous):

R=4 and D=2(4)=8